OpenStudy (anonymous):

Will give medal to best answer! The expression cos^-1 (3/5) has an infinite number of values. True or False?? I think it's false.

OpenStudy (anonymous):

@divu.mkr

OpenStudy (anonymous):

false

OpenStudy (mathmale):

\[\cos ^{-1}\frac{ 3 }{ 5 }\] has a solution in the first quadrant and another in the fourth quadrant, and that's just on the interval\[\left[ 0,2\pi\right]\]

OpenStudy (mathmale):

but it happens that the cosine function is periodic; therefore, its graph repeats itself every 2pi radians. Therefore, \[\cos ^{-1}\frac{ 3 }{ 5 }\] is far from being the only solution.

OpenStudy (mathmale):

Starting with the solution (angle) that's in the first quadrant, this inverse function has infinitely many solutions: \[\theta(plus.or.minus) n2\pi,\], for n={0,1,2,3,4,..... }

OpenStudy (mathmale):

and a similar expression can be written for the solution whose initial value is in Q4.

OpenStudy (mathmale):

So: What is your answer to the original question?

OpenStudy (anonymous):

well it was false at first..

OpenStudy (anonymous):

but now I'm thinking it's true?

OpenStudy (mathmale):

And now your answer is:

OpenStudy (anonymous):

true?

OpenStudy (anonymous):

no, it's false. Cosine inverse a one-to-one function.

OpenStudy (anonymous):

@ranga